Please help me solve for the unknowns (a and b):
Two numbers added together are 185. The difference between them is 57.
Here is the work that I have done so far:
a+b=185
a-b=57
Please help me. Thank you.

Answer by stanbon(75887)

a+b=185
a-b=57
-------------
Add the two equations to get:
2a = 242
a = 121
-----------
Substitute into a+b=185 to solve for "b":
121 + b = 185
b = 64
============
Cheers,
Stan H.